OpenStudy (anonymous):

Tell what you would do to isolate the variable. 3z = 5

OpenStudy (anonymous):

Since the variable, in this case z, is being multiplied by 3 you do the opposite so divide 3.

OpenStudy (tester97):

\[3z=5\] \[\frac{ 1 }{ 3 } \times 3z=5\times \frac{ 1 }{ 3 }\] \[z=1.6666666667\]